# Evaluating Named Entity Recognition: Comparative Analysis of Mono- and Multilingual Transformer Models on Brazilian Corporate Earnings Call Transcriptions

### v1

@@ Abstract
In recent years, Natural Language Processing (NLP) interest has surged due to abundant text data. Named Entity Recognition (NER) is a NLP technique for extracting information from textual documents. However, much of the existing research on NER has been centered around English-language documents, leaving a gap in the availability of datasets tailored to the financial domain in Portuguese. This study addresses the need for NER within the financial domain, focusing on Portuguese-language texts extracted from earnings call transcriptions of Brazilian banks. By curating a comprehensive dataset comprising 384 transcriptions and leveraging weak supervision techniques for annotation, we evaluate the performance of monolingual models trained on Portuguese (BERTimbau and PTT5) and multilingual models (mBERT and mT5). Notably, we introduce a novel approach that reframes the token classification task as a text generation problem, enabling fine-tuning and evaluation of T5 models. Following the fine-tuning of the models, we conduct an evaluation on the test dataset, employing performance and error metrics. Our findings reveal that BERT-based models consistently outperform T5-based models. Furthermore, while the multilingual models exhibit comparable macro F1-scores, BERTimbau demonstrates superior performance over PTT5. A manual analysis of sentences generated by PTT5 and mT5 unveils a high degree of similarity, ranging from 0.89 to 1.0, between the original and generated sentences. However, critical errors emerge as both models exhibit discrepancies, such as alterations to monetary and percentage values, underscoring the importance of accuracy and consistency in the financial domain. Despite these challenges, PTT5 and mT5 achieve impressive macro F1-scores of 98.52% and 98.85%, respectively, with our proposed approach. Furthermore, our study sheds light on notable disparities in memory and time consumption for inference across the models. For instance, BERTimbau consumes 4.5 GB of memory and 2 minutes for inference using the test dataset, whereas PTT5 requires 13.2 GB and 27 minutes. Additionally, our findings provide insights into the differing performance of BERT- and T5-based models for the NER task.
@@
Deep Learning, Financial Domain, Finance, Brazilian Banks, Stocks, Stock Exchange
@@
Compilation of 384 conference call transcriptions from Brazilian banks
Development of a Portuguese dataset with annotated sentences for NER in finance
Framing NER as text generation with T5, surpassing prior methods
Comparison of BERT and T5 models shows insights into their NER efficacy
The macro F1-score achieved by the models ranged from 98.52% to 98.99%
@@ 1 Introduction
In the investment field, various techniques are employed for stock selection in portfolios. These techniques encompass market trend analysis with Technical Analysis Siegel2014 , sentiment analysis of textual data from sources such as social media and news Carosia2021 , as well as fundamental analysis. Fundamental analysis involves scrutinizing historical financial metrics such as profit, revenue, liabilities, net earnings, cash flow, and more, alongside qualitative data found in press releases and quarterly or annual reports Prado2018 .
@@
Automating the extraction of information from these reports could significantly aid in their analysis. Named Entity Recognition (NER), a Natural Language Processing (NLP) technique, plays a crucial role in this context by identifying entities such as people, places, companies, dates, and currencies within text Li2022 . These entities are then labeled accordingly; for example, the word “London” may be identified and labeled as a location.
@@
Furthermore, NER serves as a vital information extraction technique and is often integrated with other NLP tasks such as summarization and question-answering Li2022 . Its versatility extends across diverse domains, including biomedical and financial sectors Alvarado2015 ; Francis2019 ; Song2021 ; Loukas2022 ; Zhang2023 , automatic speech recognition DelRio2021 , and legal contexts Araujo2018 . By accurately identifying and labeling entities in text, NER facilitates the automation of tasks and enhances efficiency across various fields.
@@
Building upon the importance of NER, it is essential to explore the different approaches to entity recognition. Recognizing entities in texts can be achieved through various methods Li2022 ; Song2021 . Traditional techniques rely on manually defined dictionaries and rules, which are contingent upon the quantity and quality of the defined words and rules. In contrast, machine learning techniques leverage pattern learning to identify entities in unseen texts, although they require extensive feature engineering and domain knowledge Song2021 . Alternatively, deep learning approaches have gained traction for entity recognition due to their ability to automatically uncover hidden features Li2022 ; Song2021 .
Regarding the application of NER in the financial domain, existing studies have primarily focused on generic entities such as Person, Location, Percentage, and Money Alvarado2015 ; Francis2019 ; DelRio2021 ; Loukas2022 ; Shah2023 . However, these analyses have been predominantly centered on English texts, overlooking the specific challenges and nuances present in other languages. Therefore, our study aims to bridge this gap by directing attention to earnings call transcriptions from Brazilian companies listed on the stock exchange, for which no dataset currently exists for NER tasks in Portuguese.
The transcriptions of those earnings calls encompass presentations by the company’s board, questions posed by the audience to the board, and their respective answers related to financial statements such as the balance sheet, income statement, and cash flow. These transcriptions possess intriguing characteristics due to variations between spoken language and their written transcriptions. This variability complicates automatic information extraction. For example, transcriptions may appear as “O lucro foi R\$ 900 milhões”, “O lucro foi de R\$ 900 MM” or even “O lucro foi de 900 milhões de reais” for the phrase “O lucro foi de 900 milhões de reais” (The profit was BRL 900 million.). Given these challenges, achieving proficient performance in NER tasks by language models is crucial.
Transformer-based models such as BERT and T5 have demonstrated efficacy in NER tasks Li2022 ; Francis2019 ; Song2021 ; Zhang2023 ; Wang2022instructionner ; Wang2023gptner . For our study, we selected two monolingual models pre-trained in Portuguese (BERTimbau Souza2020 and PTT5 Carmo2020 ) along with two multilingual models (mBERT Devlin2019 and mT5 Xue2021 ). Since T5 is a generative model, we adapted the token classification task to a sequence-to-sequence approach. We found that our approach yielded superior results compared to the method proposed by Wang et al. Wang2022instructionner .
Therefore, our primary objective is to evaluate the selected models using performance metrics such as Precision, Recall, and F1-score, alongside metrics adapted from the Fifth Message Understanding Conference (MUC-5) Chinchor1993-muc . MUC-5 metrics focus on the errors of the models, measuring incorrect, missed, and spurious classifications. Specifically, MUC-5 presents four metrics derived from incorrect, missing, and spurious classifications, providing insights into areas where the models may incorrectly classify a token or classify more tokens than the annotated dataset contains. Additionally, we assess the computational resources required for fine-tuning and inference. Our evaluation involves fine-tuning the models on earnings call transcriptions from ten Brazilian banks and manually verifying incorrectly labeled named entities.
@@
    Compilation of a comprehensive collection comprising 384 conference call transcriptions from ten prominent Brazilian banks. This dataset encompasses approximately 118,411 sentences and 3,082,526 tokens, facilitating robust analysis and experimentation.
@@
    Development of a dataset containing sentences annotated using weak supervision techniques. This dataset serves as a valuable resource for training and evaluating named entity recognition (NER) systems in the financial domain.
@@ 2 Background
The Transformer is a model architecture composed by an encoder-decoder structure Vaswani2017 . The encoder receives a sequence of symbols as input and maps them to a sequence of continuous representations. Based on this representation, the decoder generates an output sequence, one element at a time in an auto-regressive fashion. The Transformer structure is flexible and has been adapted for different purposes and architectures. Two examples of these adaptations are: the Bidirectional Encoder Representations from Transformers (BERT) Devlin2019 , which uses only the encoder structure, and the Text-to-Text Transfer Transformer (T5) Raffel2020 , based on the encoder-decoder structure. BERT and T5 can be pre-trained and fine-tuned to different tasks.
@@ 2.1 BERT-based models: mBERT and BERTimbau
Devlin et al. Devlin2019 introduced BERT in 2019. This model utilizes only the Transformer’s Encoder and was pre-trained using bidirectional (left and right context) representations from unlabeled text, with WordPiece as its tokenizer Devlin2019 . BERT can be fine-tuned with an additional output layer for various tasks, including question answering, language inference, and token classification Devlin2019 . BERT utilizes a fixed positional encoding scheme, which imposes a limitation on the input length.
BERT was pre-trained only in English Devlin2019 . However, the authors provided multilingual models (mBERT) on Hugging Face [^1] https://huggingface.co/: BERT-base multilingual cased and BERT-base multilingual uncased. These models were pre-trained with Wikipedia texts from 104 different languages and the authors recommend the version with 12 layers and 110M parameters.
Despite the introduction of mBERT, Souza, Nogueira and Lotufo presented BERTimbau in 2020 Souza2020 . This model, built on the BERT architecture, was pre-trained from scratch using BrWaC (Brazilian Web as Corpus) dataset. The authors released BERTimbau on Hugging Face in two versions: BERTimbau-base (12 layers and 110M parameters) and BERTimbau-large (24 layers and 335M parameters).

### v2

# Evaluating Named Entity Recognition: A Comparative Analysis of Mono- and Multilingual Transformer Models on a Novel Brazilian Corporate Earnings Call Transcripts Dataset

@@ Abstract
Since 2018, when the Transformer architecture was introduced, Natural Language Processing has gained significant momentum with pre-trained Transformer-based models that can be fine-tuned for various tasks. Most models are pre-trained on large English corpora, making them less applicable to other languages, such as Brazilian Portuguese. In our research, we identified two models pre-trained in Brazilian Portuguese (BERTimbau and PTT5) and two multilingual models (mBERT and mT5). BERTimbau and mBERT use only the Encoder module, while PTT5 and mT5 use both the Encoder and Decoder. Our study aimed to evaluate their performance on a financial Named Entity Recognition (NER) task and determine the computational requirements for fine-tuning and inference. To this end, we developed the Brazilian Financial NER (BraFiNER) dataset, comprising sentences from Brazilian banks’ earnings calls transcripts annotated using a weakly supervised approach. Additionally, we introduced a novel approach that reframes the token classification task as a text generation problem. After fine-tuning the models, we evaluated them using performance and error metrics. Our findings reveal that BERT-based models consistently outperform T5-based models. While the multilingual models exhibit comparable macro F1-scores, BERTimbau demonstrates superior performance over PTT5. In terms of error metrics, BERTimbau outperforms the other models. We also observed that PTT5 and mT5 generated sentences with changes in monetary and percentage values, highlighting the importance of accuracy and consistency in the financial domain. Our findings provide insights into the differing performance of BERT- and T5-based models for the NER task.
@@
Deep Learning , Financial Domain , Brazilian Banks , Stock Exchange
@@
Compilation of 384 conference call transcripts from Brazilian banks.
Development of the BraFiNER, a Portuguese dataset for Named Entity Recognition in finance.
Framing NER as text generation with T5, surpassing prior methods.
Comparison of BERT and T5 models shows insights into their NER efficacy.
The macro F1-score achieved by the models ranged from 98.33% to 98.99%.
@@ 1 Introduction
In the investment field, various techniques are employed for stock selection in portfolios. These techniques encompass market trend analysis with Technical Analysis [1], sentiment analysis of textual data from sources such as social media and news [2], as well as fundamental analysis. Fundamental analysis involves scrutinizing historical financial metrics such as profit, revenue, liabilities, net earnings, cash flow, and more, alongside qualitative data found in press releases and quarterly or annual reports [3].
@@
Automating the extraction of information from these reports could significantly aid in their analysis. Named Entity Recognition (NER), a Natural Language Processing (NLP) technique, plays a crucial role in this context by identifying entities such as people, places, companies, dates, and currencies within text [4]. These entities are then labeled accordingly; for example, the word “London” may be identified and labeled as a location.
@@
NER serves as a vital information extraction technique and is often integrated with other NLP tasks such as summarization and question-answering [4]. Its versatility extends across diverse domains, including biomedical and financial sectors [5, 6, 7, 8, 9, 10, 11, 12], automatic speech recognition [13], and legal contexts [14]. Regarding the application of NER in the financial domain, existing studies have primarily focused on generic entities such as Person, Location, Percentage, and Money [5, 6, 13, 8, 15, 10].
@@
Recognizing entities in texts can be achieved through various methods [4, 7]. Traditional techniques rely on manually defined dictionaries and rules, which are contingent upon the quantity and quality of the defined words and rules. In contrast, machine learning techniques leverage pattern learning to identify entities in unseen texts, although they require extensive feature engineering and domain knowledge [7]. Alternatively, deep learning approaches have gained traction for entity recognition due to their ability to automatically uncover hidden features [4, 7].
As a deep learning-based approach, the Transformer model architecture includes both Encoder and Decoder modules [16]. Since its publication in 2018, various Transformer-based models have been pre-trained or fine-tuned for downstream tasks. For example, Bidirectional Encoder Representations from Transformers (BERT) [17], Text-to-Text Transformer (T5) [18], and Generative Pre-Training (GPT) [19] are models based on the Transformer architecture, but they differ in their structure. While BERT is an Encoder-based model, T5 employs an Encoder-Decoder architecture, and GPT is a Decoder-based model. These models have demonstrated efficacy in NER tasks [4, 6, 7, 9, 20, 21, 10, 11, 12]. However, their application has been predominantly centered on English texts, overlooking the specific challenges and nuances present in other languages.
Given the importance of NER in the financial domain, it is essential to explore approaches to entity recognition in non-English texts, such as Brazilian Portuguese, using different model architectures. In this study, we aim to perform a comparative analysis of models with different Transformer-based architectures pre-trained in Portuguese. We compared the performance of BERT and T5-based models in the financial NER task, assessed the computational resources required for fine-tuning, and analyzed their errors in recognizing entities in financial texts written in Portuguese.
We encountered several challenges during this process. The first challenge was finding models with different architectures pre-trained in Portuguese. We selected two monolingual models pre-trained in Portuguese (BERTimbau [22] and PTT5 [23]) along with two multilingual models (mBERT [17] and mT5 [24]). The second challenge was related to the T5 architecture. Since T5 is a generative model, we had to adapt the token classification task as used in BERT to a sequence-to-sequence approach. The third challenge was the lack of a suitable dataset. We did not find a dataset based on financial texts from companies listed on the Brazilian stock exchange annotated for named entity recognition. Therefore, we prepared the Brazilian Financial NER Dataset (BraFiNER) using 384 earnings call transcripts from ten Brazilian banks and used it for model fine-tuning and evaluation.
@@
    Compilation of a comprehensive collection comprising 384 conference call transcripts from ten prominent Brazilian banks. This dataset encompasses approximately 118,411 sentences and 3,082,526 tokens, facilitating robust analysis and experimentation.
@@
    Development of the Brazilian Financial NER Dataset (BraFiNER) containing sentences annotated using weakly supervised techniques. This dataset serves as a valuable resource for training and evaluating named entity recognition (NER) systems in the financial domain.
@@
2 Related Work
Named Entity Recognition (NER) has been extensively studied over the years using various datasets, label sets, and strategies [4, 7]. The following subsections present related works on the pre-training of Transformer-based models for the financial domain, the fine-tuning of Transformer-based models, and the development of datasets for financial NER.
2.1 Pre-training Transformer-based models for the Financial domain
The application of pre-trained models in the financial domain has been extensively studied over the years [25, 26], and various Transformer-based models have been developed [26]. Examples of these models include FinBERT [27], FinBERT PT-BR [28], and FLANG-BERT and FLANG-ELECTRA [29].
FinBERT [27] is a BERT-based model pre-trained from scratch using both general and financial English corpora. The FinBERT-PT-BR [28] model is based on BERTimbau [22], another BERT-based model, but pre-trained on Brazilian Portuguese corpora. In FinBERT-PT-BR, the authors continued the pre-training of BERTimbau by adding news from the Brazilian financial market. Shah et al. [29] released two models: FLANG-BERT, based on the BERT architecture, and FLANG-ELECTRA, based on the ELECTRA architecture. These models were pre-trained on general and financial English datasets.
In these examples, the authors pre-trained their models and fine-tuned them for downstream tasks such as Sentiment Analysis and NER. Unlike these works, we fine-tuned BERT- and T5-based models that were either pre-trained exclusively in Brazilian Portuguese (monolingual) or included Brazilian Portuguese in their corpora (multilingual) for the NER task. Besides, unlike Santos et al. [28], our dataset comprises text from earnings call transcripts for NER, while they used financial news for Sentiment Analysis.
2.2 Fine-tuning of Transformer-based models for Financial NER
Regarding the fine-tuning of Transformer-based models for the NER task, BERTimbau [22], a Portuguese pre-trained BERT-based model, achieved state-of-the-art performance on the Portuguese MiniHAREM dataset. Similarly, PTT5 [23], introduced in 2020, is a Portuguese pre-trained T5 model. After pre-training, it was fine-tuned for various tasks, including NER. In their NER experiments, the authors prefixed the instruction “Recognize Entities” to sentences, as in “Recognize Entities: John lives in New York”. They expected an output following the pattern: “John [Person] lives in [Other] New York [Location]”, where the words in brackets represent the entity types. Experiments were conducted using the Portuguese datasets HAREM and MiniHAREM, widely utilized for NER tasks, resulting in an F1-score of 82%.
In 2021, Finardi et al. [30] introduced BERTaú, a BERT-based model pre-trained from scratch using data from the Brazilian bank Itaú virtual assistant. This model was fine-tuned for various tasks including Frequent Answer Question Retrieval, Sentiment Analysis, and NER. Their NER dataset, which comprises 18,370 manually annotated examples across 16 classes including specific banking entities, yielded an F1-score of 0.877 after 5 epochs of fine-tuning. Additionally, they reported F1-scores of 0.840 for mBERT uncased and 0.853 for BERTimbau base. During inference, a V100 GPU was employed, with an average processing time of 0.01944 seconds per sample.
In 2022, Wang et al. [20] introduced the InstructionNER framework, which leverages the T5-large model to address the Named Entity Recognition (NER) task by converting it from token classification to a sequence-to-sequence (seq2seq) format. Fundamentally, the approach involves feeding the model with a sentence, an instruction, and options. The sentence represents the text in which named entities are to be identified, the instruction guides the model on what action to take, and the options denote the types of entities. The model’s output adheres to the template “entity is a/an type”. The authors employed few-shot NER for fine-tuning, utilizing five datasets distinct from the financial domain. Across their experiments, they achieved a maximum F1 value of 0.954.
In 2024, Krstev et al. [12] fine-tuned multilingual BERT-based models to perform NER and Relation Extraction on analysts stock rating news. They obtained 180,000 financial-related texts written in English and German, filtered those related to analysts ratings, and manually labeled about 1000 samples. For NER, they considered eight classes related, for example, to the analyst and company’s name, the price target, and the position of the analyst.
Similar to the work of Carmo et al. [23], Finardi et al. [30], Wang et al. [20], and Krstev et al. [12], our study involved fine-tuning BERT- and T5-based models for the Named Entity Recognition (NER) task. However, our approach diverged from that of Carmo et al. [23], who introduced PTT5—a T5 model pre-trained from scratch and subsequently fine-tuned for multiple tasks. In contrast, we fine-tuned their pre-trained model and the mT5 model specifically for domain-specific NER. Additionally, while Finardi et al. [30] pre-trained BERT from scratch and fine-tuned it for a variety of tasks, our work was focused on fine-tuning and comparing two BERT-based models pre-trained on Portuguese (BERTimbau) and multilingual (mBERT) datasets, exclusively for NER. Furthermore, our study emphasized domain-specific entities such as profit, revenue, net income, and cash flow, whereas Finardi et al.’s [30] work was centered on entities related to customer services. Finally, while Krstev et al. [12] concentrated on multilingual BERT-based models applied to news regarding analysts’ stock ratings written in English and German, with a focus on specific entities, we compared monolingual and multilingual models using different Transformer architectures for texts written in Brazilian Portuguese, targeting a broader range of entities suitable for various applications.
In contrast to the approach of Wang et al. [20], who employed the large version of T5, we fine-tuned two T5-based models with fewer parameters pre-trained on Portuguese (PTT5 base version) and multilingual (mT5 small version) datasets. Similarly, comparable to the methodologies of Carmo et al. [23] and Wang et al. [20], we reformulated the NER task from token classification to seq2seq. However, our approach diverges from theirs in that we fine-tuned the models using a sentence as input and generated the same sentence as output, with entities and their labels annotated within the sentence.
2.3 Dataset development for Financial NER
In 2015, Alvarado, Verspoor, and Baldwin [5] constructed a NER dataset utilizing the CoNLL-2003 dataset available in the literature, in addition to a supplementary dataset created by them, which comprised domain-specific English texts. This supplementary dataset included texts extracted from financial agreements, sourced from 8 documents (totaling 54,256 words) randomly selected for manual annotation. They annotated four types of entities provided in the CoNLL-2003 dataset: LOCATION (LOC), ORGANIZATION (ORG), PERSON (PER), and MISCELLANEOUS (MISC). Furthermore, they automatically annotated all instances of “lender” and “borrower” as PERSON (PER). Experiments were conducted using various combinations of training and testing datasets, employing Conditional Random Fields (CRFs) for annotation. The highest F-score (82.7%) was attained when the CRF was trained and tested with the domain-specific dataset.
In 2019, Francis, Van Landeghem, and Moens [6] developed a dataset using 3000 invoices from service providers such as insurance, telecommunications, banking, and tax companies. They employed optical character recognition (OCR) to extract English texts related to International Bank Account Number (IBAN) of the beneficiary, invoice number, invoice and due dates, total inclusive amount, total exclusive amount, total Value Added Tax (VAT) amount, structured communication, company name, company address, and other related fields required for invoice processing. The NER task was performed using an architecture consisting of an input layer, two stacked Bidirectional Long Short-Term Memory (Bi-LSTM), and a CRF layer. Due to variances in entity structure between the gold dataset [^1] A gold dataset is considered to be of high quality and accuracy due to the meticulous manual annotation process conducted by human annotators. and the extracted entities, the authors evaluated their strategy manually.
Two years later, in 2021, Del Rio et al. [13] compiled a dataset using English transcripts from 44 earnings calls across nine sectors, recorded in 2020. They meticulously annotated 24 types of entities, including PERSON, ORG, LOC, PRODUCT, EVENT, WORK_OF_ART, LAW, LANGUAGE, DATE, TIME, PERCENT, MONEY, QUANTITY, ORDINAL, CARDINAL, ABBREVIATION, CONTRACTION, ALPHANUMERIC, YEAR, and WEBSITE. The annotation proceeded in three steps: 1) Their custom tool was used to annotate tokens requiring normalization, such as abbreviations, cardinals, ordinals, and contractions; 2) SpaCy 2.3.5222https://spacy.io/ was employed to annotate, for example, ORG and PERSON; and 3) Manual review of annotations and updates were conducted.
Subsequently, in 2022, Loukas et al. [8] introduced the FINER-139 dataset. This extensive dataset contains 1.1 million sentences tagged with gold XBRL (eXtensible Business Reporting Language) labels. Extracted from reports of publicly traded companies filed with the US Securities and Exchange Commission (SEC) over five years (2016 to 2020), it encompasses the 139 most prevalent XBRL entity types among nearly 6,000 entity types. The authors conducted experiments employing Bi-LSTMs and BERT to evaluate the dataset’s utility and applicability.
Shah et al. [15], in 2023, introduced a weakly supervised NER pipeline for financial texts, named FiNER, and created the FiNER-ORD dataset. This dataset comprises manually annotated financial news articles in English. From a pool of nearly 48,000 news articles, the authors selected 201 and annotated them using the Doccano [^3] https://doccano.github.io/doccano/ software. The annotations focused on generic entities such as PERSON, LOCATION, and ORGANIZATION. Subsequently, the annotated dataset was utilized to validate the efficacy of the developed pipeline.
Similarly, in 2023, Zhang et al. [9] conducted experiments involving large language models and the recognition of named entities in the Chinese financial domain. They constructed a dataset primarily consisting of text sourced from financial websites, company earnings releases, financial short messages on social media, and financial article summaries, totaling 7,521 sentences. Manual annotation was performed on these sentences, covering six types of generic entities: PERSON_NAME, PRODUCT_NAME, COMPANY_NAME, LOCATION, ORG_NAME, and TIME.
In line with previous research efforts, we have curated a dataset tailored for the NER task. However, our dataset comprises Portuguese text sourced from 384 earnings call transcripts of 10 banks listed on the Brazilian stock exchange. Employing weak supervision, we annotated the sentences, incorporating both generic and domain-specific labels.
@@ 3 Background
The Transformer is a model architecture composed by an encoder-decoder structure [16]. The encoder receives a sequence of symbols as input and maps them to a sequence of continuous representations. Based on this representation, the decoder generates an output sequence, one element at a time in an auto-regressive fashion. The Transformer structure is flexible and has been adapted for different purposes and architectures. Two examples of these adaptations are: the Bidirectional Encoder Representations from Transformers (BERT) [17], which uses only the encoder structure, and the Text-to-Text Transfer Transformer (T5) [18], based on the encoder-decoder structure. BERT and T5 can be pre-trained and fine-tuned to different tasks.
@@ 3.1 BERT-based models: mBERT and BERTimbau
Devlin et al. [17] introduced BERT in 2019. This model utilizes only the Transformer’s Encoder and was pre-trained using bidirectional (left and right context) representations from unlabeled text, with WordPiece as its tokenizer [17]. BERT can be fine-tuned with an additional output layer for various tasks, including question answering, language inference, and token classification [17]. BERT utilizes a fixed positional encoding scheme, which imposes a limitation on the input length.
BERT was pre-trained only in English [17]. However, the authors provided multilingual models (mBERT) on Hugging Face [^4] https://huggingface.co/: BERT-base multilingual cased and BERT-base multilingual uncased. These models were pre-trained with Wikipedia texts from 104 different languages and the authors recommend the version with 12 layers and 110M parameters.
Despite the introduction of mBERT, Souza, Nogueira and Lotufo presented BERTimbau in 2020 [22]. This model, built on the BERT architecture, was pre-trained from scratch using BrWaC (Brazilian Web as Corpus) dataset. The authors released BERTimbau on Hugging Face in two versions: BERTimbau-base (12 layers and 110M parameters) and BERTimbau-large (24 layers and 335M parameters).